# Entropy formula for surface diffeomorphisms

### v1

@@ 4.2 Decomposition of the volume growth
    In this case, $a_{i}\notin\Gamma^{2}_{i}$ and any time in the interval $(a_{i},a_{i+1}]$ is a ($L$-expanded) geometric time.
    Recall the definitions of $\varepsilon_{f^{p}}$, $B_{p}$ and $\hat{\Theta}$ introduced before the claim.
@@
    $\log\#\Omega=\log r(f,a_{i+1}-a_{i},\varepsilon_{f^{p}})\leq B_{p}+\left(a_{i+1}-a_{i}\right)\cdot\left(h_{\rm top}(f)+1\right).$
@@
    Applying Proposition 4.3 to each point $z\in\Omega$ and $n=a_{i+1}-a_{i}$, we have a family $\Theta_{z}$ of affine maps from $[0,1]$ to itself with the following properties:
@@
      $f^{k}\circ f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}\circ\theta$ is $\varepsilon_{f^{p}}$-bounded for any $\theta\in\Theta_{z}$ and $k\leq a_{i+1}-a_{i}$,
@@
    $\leq\#\left([m,a_{i+1})\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 2B_{p}+\left(a_{i+1}+1\right)\cdot A_{r}.$
    The remaining properties asserted in the claim are obvious. The term $2B_{p}$” arises from this case.
@@
    $\sum_{k=a_{j}}^{a_{i+1}-1}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)=\sum_{k=a_{j}}^{a_{i}}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)<a_{i+1}-a_{j}.$
@@
    Let $t\in[-1,1]$ be such that $\sigma\circ\theta_{i}\circ\theta(t)=x\in\Omega^{p,m,l,s}_{n,E,\gamma}$ for some $x$. By the second assumption in the claim, there are some $\theta_{j}\in\Theta_{j}$ and some affine maps $\widetilde{\theta}_{j},\widetilde{\theta}_{j+1},\cdots,\widetilde{\theta}_{i-1}$ with
@@
    $\leq\#\left([m,a_{i+1}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 2B_{p}+\left(a_{i+1}+1\right)\cdot A_{r}.$
    The remaining properties asserted in the claim follow directly from Proposition 4.1.
@@
Define
$O^{p,m,L,s}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,x\in B(z_{s},m,\frac{\varepsilon_{f}}{100\left\|Df\right\|})\text{ and }$
@@
where the union is taken over all $\gamma:=\left\{\left(\gamma^{\prime}_{i},\,\gamma^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$ such that
@@
such $\gamma$, since both $\Delta$ and $\Gamma_{l_{0}}^{2}$ have cardinality at most $n/L$. Indeed, the set $\Gamma_{l_{0}}^{2}$ records the right endpoints of geometric intervals, and $\Delta$ has the same cardinality as $\Gamma_{l_{0}}^{2}$. The number of geometric intervals (including isolated ones) is relatively small—at most n/Ln/L—because the complement of $E$ in $[0,n)$ is a disjoint union of intervals whose lengths are all greater than $L$.
@@
    we have
    $\log\#\widetilde{\Theta}_{l_{0}}$ $\leq\#\left([m,n)\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{l_{0}}}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+\frac{n}{L}\cdot 2B_{p}+n\cdot\left(A_{r}+1\right)$ (13)
@@
    $\sum_{k<a_{l_{0}}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}\quad\text{ and}\quad\sum_{k<a_{l_{0}}}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}.$
@@
$\sum_{k<a_{l_{0}}}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}=\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}.$
  • Geometric Bound: The estimation of the geometric component $\sum_{k\in[m,n)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ is rather delicate. We will show in the claim of Proposition 4.9 that this term is negligible, using a measure-theoretic argument. We retain this term in Proposition 4.7.
@@
    Recall that $n_{E}$ is the maximal geometric time in $[0,n)$ (if no such geometric time exists, we set $n_{E}=0$). We consider the times in $0,n_{E})\setminus E$  [^17] This part contains the neutral component of the trapping times $[0,m)$ and is treated here for technical convenience. The estimation of the remaining, geometric component will be carried out in Trapping Bound below.. We decompose $0,n_{E})\setminus E$ into maximal disjoint intervals:
@@
    We note that $d_{i}\in E$ by maximality and consequently, for any $x\in O^{p,m,l,s}_{n,E,\beta}$,
@@
    $\sum_{k\in[c_{i},d_{i})}\beta^{\prime\prime}_{k}\geq-10(d_{i}-c_{i})+\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
@@
    $\sum_{k\in[0,n_{E})\setminus E}\beta^{\prime\prime}_{k}\geq-10\#\left([0,n_{E})\setminus E\right)+\sum_{k\in[0,n_{E})\setminus E}\frac{\beta^{\prime}_{k}}{r}.$
    Hence, inserting the above lower bound on $\beta^{\prime\prime}_{k}$ into the inequality below,
@@
    $=\left(n_{E}-\#E\right)\cdot\left(\frac{10}{r-1}+\frac{\log\left\|Df\right\|}{r}\right).$
  • Trapping Bound: The estimate of $\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ for all $k\in[0,m)\setminus E$ has already been addressed in the Neutral bound above. It remains to treat the case where $k\in[0,m)\cap E$. We decompose $0,m)\cap E$ into maximal disjoint intervals:
@@
    where $m_{E},m-1]$ denotes the last interval if $m-1,m\in E$ (which means that the geometric interval is truncated at m1m-1). For each interval $c_{i},d_{i})$, by an almost identical argument as in the neutral bound above, we have
    $\sum_{k\in[c_{i},d_{i})}\beta^{\prime\prime}_{k}\geq-10(d_{i}-c_{i})+\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
@@
    $\sum_{i}\sum_{k\in[c_{i},d_{i})}\beta^{\prime\prime}_{k}\geq-10m+\sum_{i}\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
@@
    $\sum_{k\in[m_{E},m-1]}\beta^{\prime\prime}_{k}\geq-10m+\sum_{k\in[m_{E},m-1]}\frac{\beta^{\prime}_{k}}{r}-L\cdot\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$ (14)
    Hence, inserting the above lower bound on $\beta^{\prime\prime}_{k}$ into the inequality below,
@@
    $\sum_{k\in[n_{E},n)}\beta^{\prime\prime}_{k}-\frac{\beta^{\prime}_{k}}{r}\leq 2n+10\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|\right).$ (15)
@@
    $\Omega_{\theta}=\Big\{t\in[0,1]:$ $\left\lceil\log\left\|Df_{f^{k}\circ\sigma\circ\theta(t)}\right\|\right\rceil=\beta^{\prime}_{k},$
    $\left\lceil\log\left\|Df_{f^{k}\circ\sigma\circ\theta(t)}\left(\sigma^{\prime}(\theta(t)\right)\right\|\right\rceil=\beta^{\prime\prime}_{k},\quad k\in(n_{E},n)\Big\}$
    where $\sigma^{\prime}(s)$ denotes the unit tangent vector at $\sigma(s)$. Note that our focus is solely on the restriction $\sigma\circ\theta|_{\Omega_{\theta}}$. By (15), we have
@@
    Note that the term $10\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|\right)$” in the first inequality is introduced to compensate for the mismatch at time $n_{E}$.
@@
$\log{\rm Vol}\left(f^{n}\left(O^{p,m,l,s}_{n,E,\beta}\right)\right)\leq\log\#\widetilde{\Theta}_{l_{0}}+\left(n-n_{E}\right)\cdot\frac{\log\left\|Df\right\|}{r}+2n+20\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|\right).$
@@
$\quad+\frac{n}{L}\cdot 2B_{p}+n\cdot\left(A_{r}+4\right)+\frac{30n}{r-1}$
@@
$C_{n,L}:=\frac{n}{L}\cdot 2B_{p}+n\cdot\left(A_{r}+4\right)+\frac{30n}{r-1}+\left(\frac{3n}{L}\cdot\frac{1}{r-1}+\frac{L}{r-1}+\frac{2n}{L}+20\right)\cdot\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)$
@@
for some Borel measures (not necessarily probability measures) $\hat{\mu}^{L}_{p},\hat{\mu}_{p}$ and $\hat{\mu}$. Then $\hat{\mu}_{p}$ and $\hat{\mu}$ are $\hat{f}-$invariant. Moreover, the projection $\mu$ of $\hat{\mu}$ carries at least one positive Lyapunov exponent (assuming $\mu$ is not zero measure) and $\hat{\mu}$ is an unstable lift of $\mu$.
@@
Next we show that $\mu$ has at least one positive Lyapunov exponent. We first note the following two facts:
@@
Note that in the definition of $\Omega^{L}$ above, we use the larger function $\rho$ rather than $\rho^{\prime}$ for simplicity (since $\rho\geq\rho^{\prime}$). Note that the number of indices $i$ satisfying $i\in E_{n}^{L}$ and $i+1\notin E_{n}^{L}$ is at most $\frac{n}{L}$, since each such $i$ corresponds to the right endpoint of a geometric interval and the complement of $E_{n}^{L}$ consists of intervals of length at least $L$. As a consequence, for any $n,p\in\mathbb{N}$, we have
@@
$\limsup_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}\left(\hat{x}\right)\right)\geq 1>0.$
This implies the projection $\mu$ of $\hat{\mu}$ has at least one positive Lyapunov exponent and $\hat{\mu}$ is an unstable lift of $\mu$. ∎
The following proposition provides an upper bound for the volume growth of a sequence of curves, already exhibiting the convex structure that interpolates between topological entropy and the Yomdin term. It can be viewed as a preliminary form of the convex upper bound in Theorem 2.1, where the parameter $\alpha$ arises from the relative proportions of different dynamical time regimes.
@@
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$ and let $\sigma_{n}$ be a sequence of $\varepsilon_{f}$-bounded $C^{r}$ curves. There is a number $\alpha\in[0,1]$such that
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\alpha\cdot h_{\rm top}(f)+\left(1-\alpha\right)\cdot\frac{\log\left\|Df\right\|}{r}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r}$
@@
Proof.
Recall that the periodic sources are denoted by $z_{1},z_{2},z_{3},\cdots$.
@@
$t_{n}(x):=\max\left\{0<k\leq n:~\exists 1\leq s\leq n,\text{ s.t. }x\in B\left(z_{s},k,\frac{\varepsilon_{f}}{100\left\|Df\right\|}\right)\right\}$
@@
$x\notin B\left(z_{s},\frac{\varepsilon_{f}}{100\left\|Df\right\|}\right).$
@@
    where, as previously defined, E(x^)E($\hat${x}) is set of geometric times of $\hat{x}$. Let $\mathcal{E}_{n}$ be the partition of $\sigma^{m_{n}}_{n}$ with such elements $\Lambda_{E_{n}}$ for all possible $E_{n}$. Note that for the cardinality, we have
@@
    • For any $L\in\mathbb{N}$, by definitions,
      $\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\subset O^{p,m_{n},L}_{n,E^{L}_{n},\gamma}$
@@
      $\beta^{\prime}_{i}:=\gamma^{\prime}_{i},\,\,\beta^{\prime\prime}_{i}:=\hat{\gamma}^{\prime}_{i},\,\,i\in[m_{n},n)\cap E_{n}^{L}\,\,\text{ and }\,\,\beta^{\prime}_{i}:=\gamma^{\prime\prime}_{i},\,\,\beta^{\prime\prime}_{i}:=\hat{\gamma}^{\prime\prime}_{i},\,\,i\in[0,n)\setminus\left([m_{n},n)\cap E_{n}^{L}\right).$
@@
${\rm Vol}\left(f^{n}\left(\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\right)\right)\geq\max_{\widetilde{\gamma}}\,{\rm Vol}\left(f^{n}\left(\Lambda^{p,m_{n}}_{n,E_{n},\widetilde{\gamma}}\right)\right).$
@@
By Proposition 4.7, for any $p,L\in\mathbb{N}$, recalling that $\gamma$ is determined by $\Lambda_{E_{n}}$ and $p$, and that $\beta$ is determined by $\gamma$,
@@
  • $\mu$ has at least one positive Lyapunov exponent,

### v2

@@ 4.2 Decomposition of the volume growth
    In this case, $a_{i}\notin\Gamma^{2}_{i}$ and $(a_{i},a_{i+1}]\subset E$.
    Recall the definitions of $\varepsilon_{f^{p}}$, $B_{p}$ and $\hat{\Theta}$ ($\hat{\Theta}\leq B_{p}$) introduced before the claim.
@@
    $\log\#\Omega=\log r(f,a_{i+1}-a_{i},\varepsilon_{f^{p}})\leq B_{p}+\left(a_{i+1}-a_{i}\right)\cdot\left(h_{\rm top}(f)+A_{r}\right).$
@@
    Applying Proposition 4.3 to each point $z\in\Omega$ with $n=a_{i+1}-a_{i}$, we have a family $\Theta_{z}$ of affine maps from $[0,1]$ to itself with the following properties:
@@
      $f^{k}\circ f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}\circ\theta$ is $\varepsilon_{f^{p}}$-bounded for any $\theta\in\Theta_{z}$ and $0\leq k\leq a_{i+1}-a_{i}$,
@@
    $\leq\#\left([m,a_{i+1})\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 3B_{p}+a_{i+1}\cdot 2A_{r}.$
    The remaining properties asserted in the claim are obvious. The terms $3B_{p},2A_{r}$” arise from this case.
@@
    $\sum_{k=a_{j}}^{a_{i+1}-1}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)=\sum_{k=a_{j}}^{a_{i}}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)=\sum_{k=j}^{i}\rho^{\prime}\left(\hat{f}^{a_{k}}(\hat{x})\right)<a_{i+1}-a_{j}=i+1-j.$
@@
    Let $t\in[0,1]$ be such that $\sigma\circ\theta_{i}\circ\theta(t)=x\in\Omega^{p,m,L,s}_{n,E,\gamma}$ for some $x$. By the second assumption in the claim, there are some $\theta_{j}\in\Theta_{j}$ and some affine maps $\widetilde{\theta}_{j},\widetilde{\theta}_{j+1},\cdots,\widetilde{\theta}_{i-1}$ with
@@
    $\leq\#\left([m,a_{i+1}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 3B_{p}+a_{i+1}\cdot 2A_{r}.$
    The remaining properties asserted in the claim are obvious.
@@
We define a subset of $O^{p,m,L}_{n,E,\beta}$ associated with a fixed periodic source $z_{s}$ by [^30] In the case $m=0$, we define $O^{p,0,L,s}_{n,E,\gamma}$ by replacing the condition $x\in B(z_{s},m,r_{z_{s}})$” in $O^{p,m,L,s}_{n,E,\gamma}$ with $x\notin B(z_{s},r_{z_{s}})$”, as before.
$O^{p,m,L,s}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,x\in B(z_{s},m,r_{z_{s}})\text{ and }$
@@
where the union is taken over all possible $\gamma:=\left\{\left(\gamma^{\prime}_{i},\,\gamma^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$ such that
@@
such $\gamma$, since both $\Delta$ and $\Gamma_{l_{0}}^{2}$ have cardinality at most $n/L$. Indeed, the set $\Gamma_{l_{0}}^{2}$ records the right endpoints of geometric intervals, and $\Delta$ has the same cardinality as $\Gamma_{l_{0}}^{2}$. The number of geometric intervals (including isolated ones) is relatively small-at most $n/L$-because the complement of $E$ in $[0,n)$ is a disjoint union of intervals whose lengths are all larger than $L$.
@@
    recalling $a_{l_{0}}=n_{E}$,
    $\log\#\widetilde{\Theta}_{l_{0}}$ $\leq\#\left([m,n)\cap E\right)\cdot h_{\rm top}(f)+\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+\frac{n}{L}\cdot 3B_{p}+n\cdot 2A_{r}$ (16)
@@
    $\sum_{k\in[0,n_{E})}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}\quad\text{ and}\quad\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}.$
@@
$\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}.$
  • Geometric Bound:
    The estimation of the geometric component $\sum_{k\in[m,n)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ is rather delicate. We will show in Proposition 4.10 that this term is negligible, using a measure-theoretic argument. We retain this term in Proposition 4.8.
@@
    Recall that $n_{E}$ is the maximal geometric time in $[0,n)$ (if no such geometric time exists, we set $n_{E}=0$). We consider the times in $0,n_{E})\setminus E$  [^31] This part contains the neutral component (i.e., the non-geometric times $0,m)\setminus E$) of the trapping times $[0,m)$ and is treated here for technical convenience. The estimation of the remaining, geometric component (i.e., $0,m)\cap E$) will be carried out in Trapping Bound below.. We decompose $0,n_{E})\setminus E$ into maximal disjoint intervals:
@@
    We note that $d_{i}$ must be a geometric time by maximality and consequently, for any $x\in O^{p,m,L,s}_{n,E,\beta}$,
@@
    $\sum_{k\in[c_{i},d_{i})}-\beta^{\prime\prime}_{k}\leq 10(d_{i}-c_{i})-\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
@@
    $\sum_{k\in[0,n_{E})\setminus E}-\beta^{\prime\prime}_{k}\leq 10\#\left([0,n_{E})\setminus E\right)-\sum_{k\in[0,n_{E})\setminus E}\frac{\beta^{\prime}_{k}}{r}.$
    Hence,
@@
    $=\#\left([0,n_{E})\setminus E\right)\cdot\left(\frac{10}{r-1}+\frac{\log\left\|Df\right\|}{r}\right).$
  • Trapping Bound:
    The estimate of $\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ for all $k\in[0,m)\setminus E$ has already been addressed in the Neutral Bound above. It remains to treat the case where $k\in[0,m)\cap E$. We decompose $0,m)\cap E$ into maximal disjoint intervals:
@@
    where $m_{E},m-1]$ denotes the last interval if $m-1,m\in E$ (which means that the geometric interval is truncated at m1m-1). For each interval $c_{i},d_{i})$, by an almost identical argument as in the Neutral Bound, we have
    $\sum_{k\in[c_{i},d_{i})}-\beta^{\prime\prime}_{k}\leq 10(d_{i}-c_{i})-\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
@@
    $\sum_{i}\sum_{k\in[c_{i},d_{i})}-\beta^{\prime\prime}_{k}\leq 10m-\sum_{i}\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
@@
    $\sum_{k\in[m_{E},m-1]}-\beta^{\prime\prime}_{k}\leq 10m-\sum_{k\in[m_{E},m-1]}\frac{\beta^{\prime}_{k}}{r}+L\cdot\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$ (17)
    Hence,
@@
    $\sum_{k\in[n_{E},n)}\beta^{\prime\prime}_{k}-\frac{\beta^{\prime}_{k}}{r}\leq 2n+10\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$ (18)
@@
    $\Omega_{\theta}=\Big\{t\in[0,1]:$ $x=\sigma\circ\theta(t),\quad\left\lceil\log\left\|Df_{f^{k}(x)}\right\|\right\rceil=\beta^{\prime}_{k},$
    $\left\lceil\log\left\|Df\left(\hat{f}^{k}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{k},\quad k\in(n_{E},n)\Big\}.$
    Note that our focus is solely on the restriction $\sigma\circ\theta|_{\Omega_{\theta}}$. By (18), we have
@@
    Note that the term $10\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)$” in the first inequality is introduced to compensate for the mismatch at time $n_{E}$.
@@
$\log{\rm Vol}\left(f^{n}\left(O^{p,m,L,s}_{n,E,\beta}\right)\right)\leq\log\#\widetilde{\Theta}_{l_{0}}+\left(n-n_{E}\right)\cdot\frac{\log\left\|Df\right\|}{r}+2n+20\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$
@@
$\quad+\frac{n}{L}\cdot 3B_{p}+n\cdot\left(2A_{r}+3\right)+\frac{30n}{r-1}$
@@
$C_{n,L}:=\frac{n}{L}\cdot 3B_{p}+n\cdot\left(2A_{r}+3\right)+\frac{30n}{r-1}+\left(\frac{3n}{L}\cdot\frac{1}{r-1}+\frac{L}{r-1}+\frac{2n}{L}+20\right)\cdot\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)$
@@
Recall the geometric times in Definition 11 and the maximal Lyapunov exponent of $x$:
$\lambda^{+}(f,x):=\limsup_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|.$
@@
for some Borel measures (not necessarily probability measures) $\hat{\mu}^{L}_{p},\hat{\mu}_{p}$ and $\hat{\mu}$. Then $\hat{\mu}_{p}$ and $\hat{\mu}$ are $\hat{f}-$invariant. Moreover, for the projection $\mu$-almost every $x$, $\lambda^{+}(f,x)>0$ (assuming $\mu$ is not zero measure) and $\hat{\mu}$ is an unstable lift of $\mu$.
@@
Next we show that for $\mu$-almost every $x$, $\lambda^{+}(f,x)>0$. We first note the following two facts:
@@
Note that in the definition of $\Omega^{L}$ above, we use the larger function $\rho$ rather than $\rho^{\prime}$ for simplicity (since $\rho\geq\rho^{\prime}$). Note that the number of indices $i$ satisfying $i\in E_{n}^{L}$ and $i+1\notin E_{n}^{L}$ is at most $\frac{n}{L}$, since each such $i$ corresponds to the right endpoint of a maximal geometric interval and the complement of $E_{n}^{L}$ consists of intervals of length at least $L$. As a consequence, for any $n,p\in\mathbb{N}$, we have
@@
$\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}\left(\hat{x}\right)\right)=\limsup_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}\left(\hat{x}\right)\right)\geq 1>0.$
This implies for the projection $\mu$-almost every $x$, $\lambda^{+}(f,x)>0$ and $\hat{\mu}$ is an unstable lift of $\mu$ (see Section 3.2). ∎
The following proposition provides an upper bound for the volume growth of a sequence of curves, which can be viewed as a prototype of Theorem 2.1
@@
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$ and let $\sigma_{n}$ be a sequence of $\varepsilon_{f}$-bounded $C^{r}$ curves. We have
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\log\left\|Df\right\|}{r}\right\}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r}$
@@
Remark.
We prove Proposition 4.10 by reformulating it in a convex combination form. More precisely, it suffices to show that there exists $\alpha\in[0,1]$ such that
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\alpha\cdot h_{\rm top}(f)+\left(1-\alpha\right)\cdot\frac{\log\left\|Df\right\|}{r}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r}.$
The parameter $\alpha$ arises from the relative proportions of different dynamical time regimes that interpolate between topological entropy and the Yomdin term. One can see from the proof that $\alpha$ corresponds, for some point $x$, to the proportion of geometric times (see Definition 11) after excluding the times spent near periodic sources (which are referred to as trapping times in the proof of Proposition 4.10).
Proof of Proposition 4.10.
Recall that the periodic sources are denoted by $z_{1},z_{2},z_{3},\cdots$ with the corresponding radii $r_{z_{i}}$ (See Definition 4.7).
@@
$t_{n}(x):=\max\left\{0<k\leq n:~\exists 1\leq s\leq n,\text{ s.t. }x\in B\left(z_{s},k,r_{z_{s}}\right)\right\}$
@@
$x\notin B\left(z_{s},r_{z_{s}}\right).$
@@
    where, as previously defined, E(x^)E($\hat${x}) is the set of geometric times of $\hat{x}$. Let $\mathcal{E}_{n}$ be the partition of $\sigma^{m_{n}}_{n}$ with such elements $\Lambda_{E_{n}}$ for all possible $E_{n}$. Note that for the cardinality, we have
@@
    • For any $L\in\mathbb{N}$, by definitions (see the definition of $O^{p,m_{n},L}_{n,E^{L}_{n},\beta}$ in front of Proposition 4.8),
      $\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\subset O^{p,m_{n},L}_{n,E^{L}_{n},\beta}$
@@
      $\beta^{\prime}_{i}:=\gamma^{\prime}_{i},\,\,\beta^{\prime\prime}_{i}:=\hat{\gamma}^{\prime}_{i},\,\,i\in[m_{n},n)\cap E_{n}^{L}\,\,\text{ and }\,\,\beta^{\prime}_{i}:=\gamma^{\prime\prime}_{i},\,\,\beta^{\prime\prime}_{i}:=\hat{\gamma}^{\prime\prime}_{i},\,\,i\in[0,n)\setminus\left([m_{n},n)\cap E_{n}^{L}\right).$ (19)
@@
${\rm Vol}\left(f^{n}\left(\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\right)\right)=\max_{\widetilde{\gamma}}\,{\rm Vol}\left(f^{n}\left(\Lambda^{p,m_{n}}_{n,E_{n},\widetilde{\gamma}}\right)\right).$
@@
By Proposition 4.8, for any $p,L\in\mathbb{N}$, recalling that $\gamma$ is determined by $\Lambda_{E_{n}}$ and $p$, and that $\beta$ is determined by $\gamma$ as (19),
@@
Consequently, for any $p\in\mathbb{N}$,
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\log\left\|Df\right\|}{r}\right\}+\limsup_{L\to+\infty}\limsup_{n\to+\infty}\gamma_{n,L,p}.$
@@
  • $\mu$-almost every $x$ has at least one positive Lyapunov exponent (i.e., $\lambda^{+}(f,x)>0$),